# This work was supported by the Center for Identification Technology Research and the National Science Foundation, under Grants 2413228 and 2501916

### v1

@@ I Introduction
Doctoring images and videos convincingly was traditionally an expensive and time-consuming process that required a high level of skill; this is no longer the case. Advancements in the field of deep learning have given people the tools to easily and convincingly create and manipulate images and videos. While these tools have positive applications, such as social media face filters or the creation of virtual reality avatars, they also have highly malicious applications, namely deepfakes. Broadly speaking, deepfakes are images or videos in which the face of the original actor has either been reanimated or replaced with a different person’s face. Deepfakes have been increasingly used in malicious ways, including, but not limited to: spreading political misinformation[1], generating celebrity pornography[2], and conducting various scams[3], [4].
In the past, one of the largest limitations faced by deepfakes was the inability to generate them in real-time while remaining convincing [5], [6]; this is no longer the case, as recent advancements now allow them to be generated in real-time while remaining convincing [6]. With the removal of this barrier, a new concerning trend has emerged, deepfake phishing in video calls [4] [3]. This attack works very similarly to classical phishing, but instead of the scam being conducted via email or phone, the attacker convinces the victim to join a video call where they will be waiting, impersonating a trusted individual. This kind of attack has already been used to devastating effect, such as when an employee of a financial firm was tricked into sending 25 million dollars to scammers after they used this technology to impersonate the company’s chief financial officer [3]. This new form of attack is not limited to theft and has been used in political espionage, such as when an attacker impersonated a Ukrainian diplomat during a meeting with a high-ranking US senator in an attempt to gain access to classified information [4].
In this paper, we propose a proof-of-concept pipeline capable of running in real-time, specifically designed to counter deepfake phishing occurring in one-on-one video calls using point-of-gaze (PoG) tracking. The motivation behind this method is that a person‘s gaze contains rich biometric information that is used as a subtle non-verbal communicator during conversations [7], [8], [9], which we hypothesize that deepfakes will be unable to mimic. This paper will cover the creation and collection of our custom dataset, the deepfake generation, the selection of the gaze estimation method, and an in-depth investigation into the model architecture and feature selection process.
@@ II Related Work
Deepfake detection methods come in a wide variety, but tend to fall into one of two categories: error-based or biometric-based [10]. Biometric-based detection methods are often more explainable than error-based detection methods, as they rely on the extraction and analysis of intermediate biometric features that can easily be pointed out and conceptualized. These biometric features can be anything from blood flow/blush [11] to emotions [12]. On the other hand, you have error-based methods which do not have an intermediate preprocessing step and often take the entire face as input [13], [14], [15], and rely on pixel-level errors and inconsistencies. This reliance on pixel-level inconsistencies, while effective, can result in dubious explainability, as there is no easily human-interpretable feature that can be singled out as the cause.
@@ II-A Detectors using Gaze-based Biometrics
Demir et al. [16] proposes a detector that relies on explainable engineered features. Their detector extracts three different features: visual, geometric, and metric. The visual and geometric features are used to create a signal that represents them temporally, while all three are combined into a spectrogram to represent them within the frequency domain. The feature most relevant to this work is geometric, which is where the eyes actually look in space; they extract this feature using OpenFace [17], an appearance-based gaze estimation method [18]. From the geometric feature, the researchers observe that in fake videos, the left and right eyes do not move in conjunction, and the gaze transitions abruptly from frame to frame. Peng et al. [19] proposes a deepfake detection method that takes a whole image as input. The image is fed into three different feature extraction models: texture, attribute, and gaze analysis. The gaze analysis model is appearance-based and is first trained on the Gaze360 [20] dataset, which is then fine-tuned on the downstream task of forgery detection. The researchers make several observations, most relevant to this paper are their findings that the gaze vectors’ pitch and yaw are more spread out over short periods of time in deepfakes than in genuine videos. Li et al.[21] proposes a detection method using a support vector machine for classification and a model-based gaze estimation method. The proposed detector relies on two observed characteristics of eyes in deepfakes. The first is that the eyes of deepfakes are more stationary than genuine eyes, and when there is movement, it is irregular. The second observation, similar to Demir et al. [16], is that the eyes of deepfakes do not move in conjunction with one another; they refer to this as binocular coordination linkage. He et al. [13] combines an error-based detection method with 3D gaze biometrics, extracted through an appearance-based model. They use spatial inconsistencies from the gaze vectors to regularize the error-based model via mean square error and leaky feature fusion.
@@ III-A Collection Process
The dataset was collected with IRB approval, and subjects were found via various campus announcements made to the undergraduate and graduate student body; the only restriction on participation was that the participant not be legally blind. Each data collection session involved two people: a member of the research team and a random volunteer. Before the two participants started their conversation, they would both perform a calibration in which they were instructed to follow a red dot as it was displayed in various locations on the screen. After calibration, the lab member would initiate the conversation with the volunteer, which was free-form in nature, as we were not concerned with the topic of conversation, only that it was natural. The lab member participating in the conversation would ensure minimal downtime by reading another question from a list of icebreakers [22] when the conversation died down. Each session would go for at least ten minutes and would only be ended by the proctor once the conversation had reached a natural conclusion. Because of this, most of the collection are about 11 minutes in length and in one case continued naturally for 18 minutes. On occasion, the connection would be dropped or another issue would occur, causing the two participants to lose connection before the 10-minute mark. In cases where this happened, they would start again from where they left off and go for the remaining time, and the erroneous portions would be removed in post.
The video data was collected at a resolution of 1920x1080 at roughly 30 fps using a Logitech Brio 500, a commercially available monocular webcam. The audio recording and communications were done with a Logitech H390 headset.
@@ III-B Preprocessing
The first preprocessing step was to align the three modalities: audio, video, and displayed video. Where audio is the conversation between the volunteer and the team member, video is the recording of the volunteer’s face as they converse, and displayed video is what the volunteer is looking at during the conversation. After aligning the three modalities, deepfakes were generated for the videos. Next, speaker diarization was performed using combination of pyannote [23], [24], [25] and manual annotation to classify each frame into one of the following four speaking context information categories: volunteer speaking, team member speaking, both speaking, or neither speaking. Then the following intermediate data was calculated and estimated: the PoG for the deepfake and genuine videos, along with facial landmarks from the displayed video, obtained with Google’s MediaPipe facial landmark detection [26]. From the intermediate PoG and facial landmarks, a unit vector and its magnitude component were calculated from the PoG to each facial landmark per frame. These preprocessing steps allow for a significantly smaller dataset while preserving the important information. A secondary benefit is that it anonymizes the dataset by removing the participants’ faces and replacing them with landmarks [27].
@@ III-C Characteristics
In total, 51 volunteers participated in the data collection; of these, 47 were included in the final dataset. Four were lost for various reasons, such as file corruption, inability to generate convincing deepfakes, and in one case, a volunteer took a phone call during the session. The final gender split of the dataset is 77% male 23% female while this is not ideal as some research shows that men and women have different exploratory gazes [28] research on gaze during dyadic social interaction, our use case, does not find any evidence of differences during conversation [7], [8], [9] so we believe that the effect of this imbalance will be negligible if not entirely mute. An in-depth breakdown of the content of the dataset can be seen in Table I.
@@ IV Deepfake Generation
Deepfakes were generated with DeepFaceLive [29] and FaceShifter [30]. The primary consideration when choosing the deepfake generation methods was the generation speed, which needed to be real-time or as close to real-time as possible, as the goal of this research is to detect real-time deepfakes. A secondary consideration was the quality of the deepfakes generated. These two considerations greatly limited the number of generation methods, as generation speed comes at the cost of quality and vice versa. Although considerable effort was invested in developing high-quality real-time generation techniques, their quality remains lower than that of their post-processing counterparts.
@@ IV-A DeepFaceLive
DeepFaceLive [29] is based on the widely used DeepFaceLabs [31] and, as such, uses much of the same pipeline. i.e., model architecture, and training method [32]. However, there are a number of important differences that precipitated its use. It works in real time and is more flexible as the impostor model can be used with the face of any impersonator, whereas DeepFaceLab requires the training of a new model for each impostor impersonator pair, which takes upwards of 6 hours. Conveniently, DeepFaceLive comes with pretrained models for 30 different celebrities, such as Jim Carrey, Tom Holland, Margot Robbie, and Ryan Reynolds. One can also train their own model using DeepFaceLive, to do this one needs at least 5000 images of the person to be impersonated under a large variety of conditions and for best results they should be collected with the same camera that the impersonator is using. This guideline of capturing the training data with the impersonator’s webcam is effectively impossible, i.e., a malicious actor would not be able to get 5000+ images of their target using a webcam without arousing suspicion; however, it is worth noting that this guideline does not seem to have a large effect. The requirement of 5000+ images, while significantly easier, would pose a challenge if the fraudster wished to impersonate someone with a less forward-facing profile, e.g., a bank manager or mid-level manager.
@@ IV-B FaceShifter
FaceShifter is a model-based deepfake generation technique that is able to perform face swaps using only a single picture of the person to be impersonated [30]. While the quality of the deepfakes generated via this method are lower at higher resolutions than DeepFaceLive, the setup needed is practically non-existent, allowing anyone with at least a single decent quality image of their face on the internet to be impersonated with ease. When we generated deepfakes with this method, we obtained slightly under real-time results, where a 30 fps 10-minute video recorded at 1920x1080 would take 11 to 12 minutes, but this could be easily remedied by lowering the output resolution, which would additionally increase the deepfake’s quality. The deepfakes were generated from 14 different celebrity faces, all of which were also available as pre-trained models for DeepFaceLive. To determine the best possible impostor for each impersonator, the cosine similarity as calculated by InsightFace [34] was used. In a few cases, a high cosine similarity did not translate to high-quality deepfakes, and in these cases, a process of trial and error was used. It is important to note that an official implementation of FaceShifter is not publicly available, and as such, an unofficial implementation [35] was used.
@@ V Point of Gaze Estimation
Estimating the PoG is a two step process. The first step is to estimate a 3D gaze vector from a monocular webcam which is in and of itself a nontrivial and open problem [36]. The second step is much simpler and involves only a ray plane intersection calculation, where the ray is the gaze vector and the plane is the screen. We investigated five different methods for obtaining the gaze vector: OpenFace [17], Falch et al. [37], MPIIGaze [38], MPIIFaceGaze [39], and ETH-XGaze [40]. Ptgaze [41] was used for the implementation of the last three methods.
OpenFace is a popular choice for estimating 3D gaze vectors but it was not designed with a transfer to 2D in mind and as per the project‘s main github contributor a conversion to 2D would not be accurate. The model based on the ETH-XGaze dataset also suffers from the same issue. On the other hand Falch and the MPII methods were designed with a transfer to 2D screen coordinates in mind. Falches’ webcam-based gaze estimation was designed specifically to transfer a 3D gaze vector to a 2D point on a computer screen in real time. The paper was able to achieve results comparable to commercial eye tracking equipment [37]. Despite these pros we found this method ineffective as it involved a calibration step which had the side effect of calibrating out the differences between the real and deepfaked gazes. We had much more success with the MPII-based models. We attribute this to the datasets similarity to our use case, which was collected “during natural everyday laptop use” [38]. Out of the two MPII-based models, the final pipeline uses MPIIFaceGaze as it has two advantages over MPIIFace. It is more accurate having a mean angle error of 4.8 degrees [39] as opposed to 6 degrees [38], which the python implementation we used also reports [42]. Secondly, it returns a single gaze vector as opposed to a gaze vector for each eye, which is helpful as it simplifies the process of finding the PoG, as we do not need to calculate where the individual gaze vectors meet in space.
@@ V-A MPIIFaceGaze
Because this method considers the entire face, it only returns a single gaze vector, originating from the centroid of six facial landmarks; see the yellow dots in Fig.  2. As to the implementation of MPIIFaceGaze we used Ptgaze [41] which is an unofficial pytorch implementation this was chosen over the official implementation because it allowed for quicker iteration and development as it was written in python as opposed to C++ [43] while remaining just as accurate reporting a mean angle error of 4.8 which [42] which translates to a average euclidean distance error of 42 millimeters [39]. An important note is that the Ptgaze [41] implementation worked using either Dlib or MediaPipe [26] for facial landmark detection, but the system took significantly longer when using Dlib where a 1 minute would take roughly 7.8 minutes to process, as opposed to MediaPipe, which was able to process a 1 minute of data in 1 minutes and 30 seconds with no discernible difference in quality, and based on the real-time requirement, MediaPipe was chosen over Dlib.

### v2

@@ I Introduction
Doctoring images and videos convincingly was traditionally an expensive and time-consuming process that required a high level of skill; this is no longer the case. Advancements in the field of deep learning have given people the tools to easily and convincingly create and manipulate images and videos. While these tools have positive applications, such as social media face filters or the creation of virtual reality avatars, they also have highly malicious applications, namely deepfakes. Broadly speaking, deepfakes are images or videos in which the face of the original actor has either been reanimated or replaced with a different person’s face. Deepfakes have been increasingly used in malicious ways, including, but not limited to: spreading political misinformation[polimisinfo], generating celebrity pornography[dfpornincident], and conducting various scams[finance_scam], [impersonation].
In the past, one of the largest limitations faced by deepfakes was the inability to generate them in real-time while remaining convincing [jpmorganAIimposter], [mittal2024gotcharealtimevideodeepfake]; this is no longer the case, as recent advancements now allow them to be generated in real-time while remaining convincing [mittal2024gotcharealtimevideodeepfake]. With the removal of this barrier, a new concerning trend has emerged, deepfake phishing in video calls [impersonation] [finance_scam]. This attack works very similarly to classical phishing, but instead of the scam being conducted via email or phone, the attacker convinces the victim to join a video call where they will be waiting, impersonating a trusted individual. This kind of attack has already been used to devastating effect, such as when an employee of a financial firm was tricked into sending 25 million dollars to scammers after they used this technology to impersonate the company’s chief financial officer [finance_scam]. This new form of attack is not limited to theft and has been used in political espionage, such as when an attacker impersonated a Ukrainian diplomat during a meeting with a high-ranking US senator in an attempt to gain access to classified information [impersonation].
In this paper, we propose a proof-of-concept pipeline capable of running in real-time, specifically designed to counter deepfake phishing occurring in one-on-one video calls using point-of-gaze (PoG) tracking. The motivation behind this method is that a person‘s gaze contains rich biometric information that is used as a subtle non-verbal communicator during conversations [dualeyetrackingconvo], [eyetrackingsetupevel], [gazesignals], which we hypothesize that deepfakes will be unable to mimic. This paper will cover the creation and collection of our custom dataset, the deepfake generation, the selection of the gaze estimation method, and an in-depth investigation into the model architecture and feature selection process.
@@ II Related Work
Deepfake detection methods come in a wide variety, but tend to fall into one of two categories: error-based or biometric-based [s22124556]. Biometric-based detection methods are often more explainable than error-based detection methods, as they rely on the extraction and analysis of intermediate biometric features that can easily be pointed out and conceptualized. These biometric features can be anything from blood flow/blush [DBLP:journals/corr/abs-1901-02212] to emotions [DBLP:journals/corr/abs-2003-06711]. On the other hand, you have error-based methods which do not have an intermediate preprocessing step and often take the entire face as input [2024106636_He], [dfdetectionRNN], [sun2021dualcontrastivelearninggeneral], and rely on pixel-level errors and inconsistencies. This reliance on pixel-level inconsistencies, while effective, can result in dubious explainability, as there is no easily human-interpretable feature that can be singled out as the cause.
@@ II-A Detectors using Gaze-based Biometrics
Demir et al. [DBLP:journals/corr/abs-2101-01165] proposes a detector that relies on explainable engineered features. Their detector extracts three different features: visual, geometric, and metric. The visual and geometric features are used to create a signal that represents them temporally, while all three are combined into a spectrogram to represent them within the frequency domain. The feature most relevant to this work is geometric, which is where the eyes actually look in space; they extract this feature using OpenFace [openface2], an appearance-based gaze estimation method [openfaceeyetracking]. From the geometric feature, the researchers observe that in fake videos, the left and right eyes do not move in conjunction, and the gaze transitions abruptly from frame to frame. Peng et al. [10478974_Peng] proposes a deepfake detection method that takes a whole image as input. The image is fed into three different feature extraction models: texture, attribute, and gaze analysis. The gaze analysis model is appearance-based and is first trained on the Gaze360 [gaze360_2019] dataset, which is then fine-tuned on the downstream task of forgery detection. The researchers make several observations, most relevant to this paper are their findings that the gaze vectors’ pitch and yaw are more spread out over short periods of time in deepfakes than in genuine videos. Li et al.[9413139_Li] proposes a detection method using a support vector machine for classification and a model-based gaze estimation method. The proposed detector relies on two observed characteristics of eyes in deepfakes. The first is that the eyes of deepfakes are more stationary than genuine eyes, and when there is movement, it is irregular. The second observation, similar to Demir et al. [DBLP:journals/corr/abs-2101-01165], is that the eyes of deepfakes do not move in conjunction with one another; they refer to this as binocular coordination linkage. He et al. [2024106636_He] combines an error-based detection method with 3D gaze biometrics, extracted through an appearance-based model. They use spatial inconsistencies from the gaze vectors to regularize the error-based model via mean square error and leaky feature fusion.
@@ III-A Collection Process
The dataset was collected with IRB approval, and subjects were found via various campus announcements made to the undergraduate and graduate student body; the only restriction on participation was that the participant not be legally blind. Each data collection session involved two people: a member of the research team and a random volunteer. Before the two participants started their conversation, they would both perform a calibration in which they were instructed to follow a red dot as it was displayed in various locations on the screen. After calibration, the lab member would initiate the conversation with the volunteer, which was free-form in nature, as we were not concerned with the topic of conversation, only that it was natural. The lab member participating in the conversation would ensure minimal downtime by reading another question from a list of icebreakers [icebreakers] when the conversation died down. Each session would go for at least ten minutes and would only be ended by the proctor once the conversation had reached a natural conclusion. Because of this, most of the collection are about 11 minutes in length and in one case continued naturally for 18 minutes. On occasion, the connection would be dropped or another issue would occur, causing the two participants to lose connection before the 10-minute mark. In cases where this happened, they would start again from where they left off and go for the remaining time, and the erroneous portions would be removed in post.
The video data was collected at a resolution of 1920x1080 at roughly 30 fps using a Logitech Brio 500, a commercially available monocular webcam. The audio recording and communications were done with a Logitech H390 headset. All of the data was collected in a laboratory where we had a large amount of control, i.e. in the lab we ensured that all of the subjects started at roughly eye level with the screen and approximately two to two and half feet from the screen, the lighting and background were also controlled and remained constant. This is not ideal but we determined it was necessary as our research into real-time point of gaze tracking using monocular webcams indicated that the technology is very sensitive to and likely would not have been usable if the aforementioned variations we controlled for were included. Additionally we believe that for the purposes of a proof of concept the controlled environment is permissible as if the system cannot work under laboratory conditions it would not be worth further investigation.
@@ III-B Preprocessing
The first preprocessing step was to align the three modalities: audio, video, and displayed video. Where audio is the conversation between the volunteer and the team member, video is the recording of the volunteer’s face as they converse, and displayed video is what the volunteer is looking at during the conversation. After aligning the three modalities, deepfakes were generated for the videos. Next, speaker diarization was performed using combination of pyannote [Bredin2020], [Bredin2021], [pyanotelink] and manual annotation to classify each frame into one of the following four speaking context information categories: volunteer speaking, team member speaking, both speaking, or neither speaking. Then the following intermediate data was calculated and estimated: the PoG for the deepfake and genuine videos, along with facial landmarks from the displayed video, obtained with Google’s MediaPipe facial landmark detection [mediapipefacelandmarks]. From the intermediate PoG and facial landmarks, a unit vector and its magnitude component were calculated from the PoG to each facial landmark per frame. These preprocessing steps allow for a significantly smaller dataset while preserving the important information. A secondary benefit is that it anonymizes the dataset by removing the participants’ faces and replacing them with landmarks [pg99-qs82-25].
@@ III-C Characteristics
In total, 51 volunteers participated in the data collection; of these, 47 were included in the final dataset. Four were lost for various reasons, such as file corruption, inability to generate convincing deepfakes, and in one case, a volunteer took a phone call during the session. The final gender split of the dataset is 77% male 23% female, which is reflective of the demographics of the pool from which the volunteers were pulled from. While this split is not ideal as some research shows that men and women have different exploratory gazes [menandwomengazeexploration] research on gaze during dyadic social interaction, our use case, does not find any evidence of differences during conversation [dualeyetrackingconvo], [eyetrackingsetupevel], [gazesignals] so we believe that the effect of this imbalance will be negligible if not entirely mute. Additionally we investigated transfer learning as a possible way to supplement our dataset, but unfortunately of the related work we found we determined that none of them were similar enough to ours to transfer from. An in-depth breakdown of the content of the dataset can be seen in Table I.
@@ IV Deepfake Generation
Deepfakes were generated with DeepFaceLive [dflive] and FaceShifter [faceshiter]. The primary consideration when choosing the deepfake generation methods was the generation speed, which needed to be real-time or as close to real-time as possible, as the goal of this research is to detect real-time deepfakes. A secondary consideration was the quality of the deepfakes generated. These two considerations greatly limited the number of generation methods, as generation speed comes at the cost of quality and vice versa. Although considerable effort was invested in developing high-quality real-time generation techniques, their quality remains lower than that of their post-processing counterparts.
@@ IV-A DeepFaceLive
DeepFaceLive [dflive] is based on the widely used DeepFaceLabs [dflabs] and, as such, uses much of the same pipeline. i.e., model architecture, and training method [perovDFlab]. However, there are a number of important differences that precipitated its use. It works in real time and is more flexible as the impostor model can be used with the face of any impersonator, whereas DeepFaceLab requires the training of a new model for each impostor impersonator pair, which takes upwards of 6 hours. Conveniently, DeepFaceLive comes with pretrained models for 30 different celebrities, such as Jim Carrey, Tom Holland, Margot Robbie, and Ryan Reynolds. One can also train their own model using DeepFaceLive, to do this one needs at least 5000 images of the person to be impersonated under a large variety of conditions and for best results they should be collected with the same camera that the impersonator is using. This guideline of capturing the training data with the impersonator’s webcam is effectively impossible, i.e., a malicious actor would not be able to get 5000+ images of their target using a webcam without arousing suspicion; however, it is worth noting that this guideline does not seem to have a large effect. The requirement of 5000+ images, while significantly easier, would pose a challenge if the fraudster wished to impersonate someone with a less forward-facing profile, e.g., a bank manager or mid-level manager.
@@ IV-B FaceShifter
FaceShifter is a model-based deepfake generation technique that is able to perform face swaps using only a single picture of the person to be impersonated [faceshiter]. While the quality of the deepfakes generated via this method are lower at higher resolutions than DeepFaceLive, the setup needed is practically non-existent, allowing anyone with at least a single decent quality image of their face on the internet to be impersonated with ease. When we generated deepfakes with this method, we obtained slightly under real-time results, where a 30 fps 10-minute video recorded at 1920x1080 would take 11 to 12 minutes, but this could be easily remedied by lowering the output resolution, which would additionally increase the deepfake’s quality. The deepfakes were generated from 14 different celebrity faces, all of which were also available as pre-trained models for DeepFaceLive. To determine the best possible impostor for each impersonator, the cosine similarity as calculated by InsightFace [insightface] was used. In a few cases, a high cosine similarity did not translate to high-quality deepfakes, and in these cases, a process of trial and error was used. It is important to note that an official implementation of FaceShifter is not publicly available, and as such, an unofficial implementation [faceshiterimplementation] was used.
@@ V Point of Gaze Estimation
Estimating the PoG is a two step process. The first step is to estimate a 3D gaze vector from a monocular webcam which is in and of itself a nontrivial and open problem [cheng2024appearancebasedgazeestimationdeep]. The second step is much simpler and involves only a ray plane intersection calculation, where the ray is the gaze vector and the plane is the screen. We investigated five different methods for obtaining the gaze vector: OpenFace [openface2], Falch et al. [Webcam-based-gaze-estimation], MPIIGaze [zhang15_cvpr], MPIIFaceGaze [zhang17_cvprw], and ETH-XGaze [zhang2020ethxgazelargescaledataset]. Ptgaze [ptgaze] was used for the implementation of the last three methods.
OpenFace is a popular choice for estimating 3D gaze vectors but it was not designed with a transfer to 2D in mind and as per the project‘s main github contributor a conversion to 2D would not be accurate. The model based on the ETH-XGaze dataset also suffers from the same issue. On the other hand Falch and the MPII methods were designed with a transfer to 2D screen coordinates in mind. Falches’ webcam-based gaze estimation was designed specifically to transfer a 3D gaze vector to a 2D point on a computer screen in real time. The paper was able to achieve results comparable to commercial eye tracking equipment [Webcam-based-gaze-estimation]. Despite these pros we found this method ineffective as it involved a calibration step which had the side effect of calibrating out the differences between the real and deepfaked gazes. We had much more success with the MPII-based models. We attribute this to the datasets similarity to our use case, which was collected “during natural everyday laptop use” [zhang15_cvpr]. Out of the two MPII-based models, the final pipeline uses MPIIFaceGaze as it has two advantages over MPIIFace. It is more accurate having a mean angle error of 4.8 degrees [zhang17_cvprw] as opposed to 6 degrees [zhang15_cvpr], which the python implementation we used also reports [pytorchmpiigaze]. Secondly, it returns a single gaze vector as opposed to a gaze vector for each eye, which is helpful as it simplifies the process of finding the PoG, as we do not need to calculate where the individual gaze vectors meet in space.
@@ V-A MPIIFaceGaze
Because this method considers the entire face, it only returns a single gaze vector, originating from the centroid of six facial landmarks; see the yellow dots in Fig.  2. As to the implementation of MPIIFaceGaze we used Ptgaze [ptgaze] which is an unofficial pytorch implementation this was chosen over the official implementation because it allowed for quicker iteration and development as it was written in python as opposed to C++ [mpiifacegaze_project] while remaining just as accurate reporting a mean angle error of 4.8 which [pytorchmpiigaze] which translates to a average euclidean distance error of 42 millimeters [zhang17_cvprw]. An important note is that the Ptgaze [ptgaze] implementation worked using either Dlib or MediaPipe [mediapipefacelandmarks] for facial landmark detection, but the system took significantly longer when using Dlib where a 1 minute would take roughly 7.8 minutes to process, as opposed to MediaPipe, which was able to process a 1 minute of data in 1 minutes and 30 seconds with no discernible difference in quality, and based on the real-time requirement, MediaPipe was chosen over Dlib.